OpenStudy (anonymous):

If the ,midponit of sr is 6,-10, and s is 0,8 , what is the point of r?

OpenStudy (campbell_st):

the mid point is the average of the x values and average of the y values so 6 = (0+ x)/2 and -10 = (8 + y)/2 so just solve x and y to find the coordinates of r

myininaya (myininaya):

\[\text{ midpoint } =(\frac{x_1+x_2}{2},\frac{y_1,y_2}{2}) \text{ where } \\ (x_1,y_1) \text{ and } (x_2,y_2) \text{ are endpoints }\] So you are given midpoint=(6,-10) and one endpoint=(0,8) So you have two equations to solve: \[\frac{0+x}{2}=6 \\ \frac{8+y}{2}=-10\]

myininaya (myininaya):

@campbell_st wins! :)

OpenStudy (campbell_st):

lol... I hate formulae like these easier to remember average of x and y

myininaya (myininaya):

yes I have told people average of a and b and they don't know what that means